Factor the following binomials.
25x^2 - 81y^2
Help me.
Thank you.

Answer by sdmmadam@yahoo.com(530)

25x^2 - 81y^2
= [(5x)^2)-(9y)^2]
= [(a)^2 - (b)^2] where a = 5x and b = 9y
= (a+b)(a-b) by formula
= (5x+9y)(5x-9y)
Answer: (25x^2 - 81y^2) = (5x+9y)(5x-9y)
Verification: Expanding the product (5x+9y)(5x-9y)
=5x(5x-9y)+9y(5x-9y) 9first term of the first bracket(along with the sign) mutiplied by the whole of the next bracket and similarly 2nd term of the ....)
=25x^2-45xy+45xy-81y^2
=25x^2+0-81y^2
=25x^2 - 81y^2